Nicholas of Derme has $200,000 invested in a Bitcoin bank that pays 38.4% annually. How long (N) will it take for his funds to quadruple?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

(4*200,000)=200,000*(1.384)^n

4=(1.384)^n

Taking log on both sides;

log 4=n*log (1.384)

n=log 4/log (1.384)

=4.27 years(Approx).